# Conformal Prediction Sets with Improved Conditional Coverage using Trust Scores

### v1

@@ 1 Introduction
Our choice of the variable $V$ is driven by an analysis of how miscoverage events are distributed across test instances in standard conformal prediction. In the classification setting, we select $V$ as a statistic that identifies instances where conformal prediction is most likely to fail—specifically when the classifier is overconfident in its incorrect predictions. Figure 1 illustrates the relationship between miscoverage rates in standard conformal prediction and the classifier’s overconfidence as measured by its reported confidence (top softmax output) and the rank of $Y_{n+1}$ in its sorted softmax probabilities. The miscoverage patterns in Figure 1 suggest that these two factors are predictive of whether conformal prediction over- or under-covers a test instance.
@@ 2.2 Dissecting miscoverage patterns in standard conformal prediction
Let Rank$(X_{n+1},Y_{n+1})$ be the position of softmax output $\hat{f}_{Y_{n+1}}(X_{n+1})$ in the decreasing sorted order of class probabilities $\hat{f}_{(1)}(X_{n+1}),\dots,\hat{f}_{(|{\mathcal{Y}}|)}(X_{n+1})$, and Conf$(X_{n+1})=\max_{y}\hat{f}_{y}(X_{n+1})$. The variables Rank$(x,y)$ and Conf$(x)$ measure a model’s overconfidence—a model is considered overconfident when it assigns a high value on its largest softmax output while the true class ranks poorly in the descending order of predicted class probabilities. In Figure 1, we visualize empirically the coverage rates of standard in different strata of Conf and Rank across various image classification datasets. As we can see, if the model’s prediction is incorrect (i.e., Rank$(X_{n+1},Y_{n+1})>1$), an increase in overconfidence leads to a reduction in coverage probability. This monotonic relationship holds consistently across all predictions that share a similar value for Rank. Additionally, we observe that for a fixed Conf, coverage probability decreases as Rank increases.
This finding suggests that, given the values of Conf and Rank for a test instance, we can determine its position within the strata in Figure 1, predict whether standard will likely over- or under-cover the true label, and adjust the conformal prediction set accordingly to achieve more balanced coverage across test instances. These miscoverage patterns are a result of the conformity score intended to approximate the oracle algorithm, and appear to hold across different datasets irrespective of the underlying data distribution. Hence, this finding is expected to generalize across datasets and likely not be alleviated by popular calibration methods such as temperature scaling as seen in Figure 1.
@@ 3.2 Algorithm
$\hat{g}_{s}=\operatorname*{arg\,min}_{g\in\mathcal{F}}\mbox{$\frac{1}{n+1}\sum_{i=1}^{n}$}\ell_{\alpha}(g(X_{i}),s_{i})+\mbox{$\frac{1}{n+1}$}\ell_{\alpha}(g(X_{n+1}),s),$ (10)
@@
For a finite-dimensional linear class =$\mathcal{F}=\{\Phi(\cdot)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$ over the basis $\Phi:\mathcal{X}\to\mathbb{R}^{d}$ibbs et al. (2023) show that we can achieve an upper bound on coverage in (9) and the exact coverage guarantee with appropriate randomization.
@@ 4 Experiments
We evaluate conditional empirically on four large-scale image classification datasets and a clinical dataset in dermatology. We propose evaluation metrics that measure the gap in coverage from the desired $1-\alpha$ level across different regions of the distribution. We also introduce a naive implementation of our proposed approach that aims to improve coverage with respect to $V$ via group-conditional coverage guarantees over non-overlapping subgroups in $V$ (Section 4.2). We evaluate standard, conditional (NAIVE), and conditional on all datasets, and show that conditional achieves the best conditional coverage performance across all settings. conditional also improves class-conditional coverage on all but one dataset.
@@ B.2 Datasets and models
Places365 [Zhou et al., 2018] contains 10 million images from 365 classes. For Places365, we train a ResNet152 model on our own (pretrained on ImageNet). We fine-tune on the train split of the original dataset and split the validation set evenly into calibration and test splits.

### v2

@@ 1 Introduction
Our choice of the variable $V$ is driven by an analysis of how miscoverage events are distributed across test instances in standard conformal prediction. In the classification setting, we select $V$ as a statistic that identifies instances where conformal prediction is most likely to fail—specifically when the classifier is overconfident in its incorrect predictions. Figure 1 (top) illustrates the relationship between miscoverage rates in standard conformal prediction and the classifier’s overconfidence as measured by its reported confidence (top softmax output) and the rank of $Y_{n+1}$ in its sorted softmax probabilities. The miscoverage patterns in Figure 1 suggest that these two factors are predictive of whether conformal prediction over- or under-covers a test instance.
@@ 2.2 Dissecting miscoverage patterns in standard conformal prediction
Let Rank$(X_{n+1},Y_{n+1})$ be the position of softmax output $\hat{f}_{Y_{n+1}}(X_{n+1})$ in the decreasing sorted order of class probabilities $\hat{f}_{(1)}(X_{n+1}),\dots,\hat{f}_{(|{\mathcal{Y}}|)}(X_{n+1})$, and Conf$(X_{n+1})=\max_{y}\hat{f}_{y}(X_{n+1})$. The variables Rank$(x,y)$ and Conf$(x)$ measure a model’s overconfidence—a model is considered overconfident when it assigns a high value on its largest softmax output while the true class ranks poorly in the descending order of predicted class probabilities. In Figure 1 (top), we visualize empirically the coverage rates of standard in different strata of Conf and Rank across various image classification datasets. As we can see, if the model’s prediction is incorrect (i.e., Rank$(X_{n+1},Y_{n+1})>1$), an increase in overconfidence leads to a reduction in coverage probability. This monotonic relationship holds consistently across all predictions that share a similar value for Rank. Additionally, we observe that for a fixed Conf, coverage probability decreases as Rank increases.
This finding suggests that, given the values of Conf and Rank for a test instance, we can determine its position within the strata in Figure 1, predict whether standard will likely over- or under-cover the true label, and adjust the conformal prediction set accordingly to achieve more balanced coverage across test instances. These miscoverage patterns are a result of the conformity score intended to approximate the oracle algorithm, and hold across different datasets irrespective of the underlying data distribution. Hence, this finding is expected to generalize across datasets and likely not be alleviated by popular calibration methods such as temperature scaling as seen in Figure 1.
@@ 3.2 Algorithm
$\hat{g}_{s}=\operatorname*{arg\,min}_{g\in\mathcal{F}}\frac{1}{n+1}\sum_{i=1}^{n}\ell_{\alpha}(g(X_{i}),s_{i})+\frac{1}{n+1}\ell_{\alpha}(g(X_{n+1}),s),$ (10)
@@
For a finite-dimensional linear class =$\mathcal{F}=\{\Phi(\cdot)^{\top}\beta:\beta\in\mathbb{R}^{d}\}$ over the basis $\Phi:\mathcal{X}\to\mathbb{R}^{d}$ibbs et al. (2023) show that we can achieve an upper bound on coverage in (9).
@@ 4 Experiments
We evaluate conditional empirically on four large-scale image classification datasets and a clinical dataset in dermatology. We propose evaluation metrics that measure the gap in coverage from the desired $1-\alpha$ level across different regions of the distribution. We also introduce a naive implementation of our proposed approach that aims to improve coverage with respect to $V$ via group-conditional coverage guarantees over non-overlapping subgroups based on $V$ (Section 4.2). We evaluate standard, conditional (NAIVE), and conditional on all datasets, and show that conditional achieves the best conditional coverage performance across all settings. conditional also improves class-conditional coverage on all but one dataset.
@@ B.2 Datasets and models
Places365 [Zhou et al., 2018] contains 10 million images from 365 classes. For Places365, we train a ResNet152 model on our own (pretrained on ImageNet). We fine-tune on the train split of the original dataset and split the validation set evenly into calibration and test splits based on the random seed.